# SymBoltz.jl: a symbolic-numeric, approximation-free and differentiable linear Einstein-Boltzmann solver

### v1

@@ 2 Code architecture and main features
SymBoltz is built somewhat differently than other Boltzmann codes. In short, it uses a symbolic-numeric abstraction interface where users enter high-level symbolic equations, and automatically compiles them to fast low-level numerical functions. It cures stiffness with implicit ODE solvers, and uses no approximation schemes to keep models simple, elegant and extensible. It is differentiable, so one can easily get accurate derivatives of any output with respect to any input. This provides rapid model prototyping, helpful abstractions and automation of typical chores when modifying other codes. SymBoltz encourages interactive usage and pursues a modular design that lets users integrate whatever they need from the package into their own applications. The end goal is to prioritize the modeler, who should be able to just write down the model equations while SymBoltz takes care of the rest.
SymBoltz is written in the Julia programming language (Bezanson et al., 2017), which attracts development of modern numerical methods and aims to resolve the two-language problem. The symbolic-numeric interface is built on the ModelingToolkit.jl (Ma et al., 2022) and Symbolics.jl (Gowda et al., 2022) packages. The compiled functions are integrated by implicit ODE solvers in DifferentialEquations.jl (Rackauckas & Nie, 2017). Everything works with automatic differentiation through ForwardDiff.jl (Revels et al., 2016), for example.
The next subsections present the design of SymBoltz in more detail, revolving around the three key features in the title. Other minor implementation details are given in appendix A. This paper is as of SymBoltz version 0.8.1. Please refer to the package documentation for definitive up-to-date information.
@@ 2.1 Symbolic-numeric interface
The core of SymBoltz is built around a symbolic-numeric interface shown in fig. 1. Variables and equations are specified in a high-level user-friendly symbolic modeling language, then ultimately compiled down to efficient low-level numerical functions that are integrated by ODE solvers. This fits the interactive and just-in-time compiled nature of Julia. A key realization is that knowledge of symbolic equations makes it possible to analyze their structure programmatically, enabling powerful transformations of the equations and automation of chores.
@@
SymBoltz turns the layout of traditional Boltzmann solvers inside-out. While other codes are built like rigid pipelines following computational stages like the background and perturbations, SymBoltz is primarily structured around physical components of the Einstein-Boltzmann system. Full cosmological models are built by joining submodels for the metric, gravitational theory, photons, baryons, dark matter, dark energy, neutrinos and other species. Each component is a chunk of related variables and internal equations, and is unaware of other components. Interactions (e.g. Compton scattering or sourcing of gravity) are equations that connect components. This structure is made possible by the symbolic interface and would slow a purely numerical code.
Separation of computational stages is secondary and done automatically (see section 2.1.3). This makes it possible to write everything related to one component in one place, instead of scattering it across separate modules for input, background, perturbations and so on. One component can be replaced by another to construct a different cosmological model without polluting a big global model. A large part of SymBoltz is thus simply devoted to building a well-organized library of such physical components.
This modular structure scales better in model space and makes it easier to assemble both extended and reduced models. While many Boltzmann solvers hardcode baryons and photons because they are fundamental to the code, SymBoltz works naturally even with pure models with non-interacting radiation, matter and dark energy. For example, this can help understand how a modified gravity theory responds to different species without complex interactions or thermodynamics cluttering the picture.
@@
Maybe this structure is unfamiliar to some. As a compromise, SymBoltz also provides an ?unstructured? version of the $\mathrm{\Lambda CDM}$ model with all equations and variables in one big system. This makes it easy-peasy to change anything in the model, and really shows the power of the symbolic interface: SymBoltz defines the full $\mathrm{\Lambda CDM}$ model (with GR, baryons, RECFAST recombination, photons, cosmological constant, massive and massless neutrinos) in just 277 lines of code, while the equivalent code that one must realistically browse in CLASS is spread over 10 files with 27721 lines! [^8] Counting input.{h,c}, background.{h,c}, thermodynamics.{h,c}, perturbations.{h,c} and wrap_recfast.{h,c} with wc -l. SymBoltz has less features than CLASS, but this is not even remotely close to making up for the $100\times$ simplification.
Let us see which helpful features this interface provides.
@@ 2.1.1 Automatic numerical code generation
$\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}\tau}=\mathbf{f}(\tau,\mathbf{u}).$ (3)
The generated code is fast and prevents users unfamiliar with the language or package from writing slow numerical code. If necessary, one can escape the standard code generation and include arbitrary numerical code, but the user is then fully responsible for its performance. This accommodates functions that cannot be written as straightforward equations, for example if one must solve a nonlinear equation for the minimum of a potential or interpolate tabulated data. The code generation deals with chores like dynamically allocating indices for each ODE state $u_{i}$, and performs optimizations such as common subexpression elimination. This is helpful as Boltzmann solvers tend to have big $\boldsymbol{f}$, and particularly with implicit ODE solvers that call $\boldsymbol{f}$ often.
@@ 2.1.2 Automatic handling of observed variables
In general, an ODE (3) admits two types of variables: unknowns $\mathbf{u}(\tau)$ are the variables that are integrated with respect to time, and observed variables are functions of the unknowns. The Einstein-Boltzmann equations are commonly formulated with many observed variables.
For example, consider the metric and gravity equations in sections A.1 and A.2 sourced by some arbitrary $\rho(\tau)$, $\delta\rho(\tau,k)$ and $\Pi(\tau,k)$. Here $a(\tau)$ and $\Phi(\tau,k)$ are the only unknown (differential) variables that the ODE solver must solve for, while one can ?observe? $z(\tau)$, $\mathscr{H}(\tau)$, $H(\tau)$, $\chi(\tau)$ and $\Psi(\tau,k)$ from the unknowns. Of course, it is always possible to eliminate all observeds by explicitly inserting them into the equations for the unknowns. But this kills readability, as observed variables are helpful and natural intermediate definitions that break up the system, and one may want to extract them from the solution as well. Furthermore, modified models can change the sets of unknown and observed variables (e.g. modified gravity can change the constraint equation for $\Psi$ into a differential equation). It is easier to compose models when variables do not have to be hardcoded as either unknown or observed.
SymBoltz reads in an entire system of equations like that defined by appendix A and automatically figures out which variables are unknown and observed. It then generates numerical code for solving the ODE for the unknowns (see section 2.1.1), and can lazily compute any observed variable or expression thereof from the solution of the unknowns.
The bottom line is that the user can trivially use and obtain any variable anywhere just by referring to it, whether it is unknown or observed. In other solvers one must look up the desired expression for observed variables and recompute them manually. This is tedious, error-prone and can tempts users to neglect parts of expressions (e.g. approximating $\Psi\approx\Phi$ by neglecting anisotropic stress).
@@ 2.1.3 Automatic stage separation and splining of unknowns
In principle, one can solve the Einstein-Boltzmann equations by integrating the entire system (i.e. background and perturbations) at the same time. However, the system can be broken down into sequential computational stages that each depend only on those before it. To alleviate stiffness in each stage, separate concerns by solving every equation only once (e.g. avoid recomputing the background for every perturbation mode) and to improve performance, it is common to solve the system stage-by-stage and spline variables from one stage as input to the next.
@@
To illustrate, again consider the general relativistic equations in section A.2 sourced by some given $\rho(\tau)$, $\delta\rho(\tau,k)$ and $\Pi(\tau,k)$. Clearly $\Phi(\tau,k)$ and $\Psi(\tau,k)$ depend on $a(\tau)$, but $a(\tau)$ does not depend on $\Phi(\tau,k)$ or $\Psi(\tau,k)$ (this just reflects the perturbative nature of the problem). One can first solve for only $a(\tau)$ in the background, then spline $a(\tau)$ and use it as input for solving for $\Phi(\tau,k)$ and $\Psi(\tau,k)$ in the perturbations, instead of solving for all three together and repeatedly solve for $a(\tau)$ for every $k$.
@@
SymBoltz uses the same stage separation strategy, but automates and strengthens it with its knowledge of the symbolic equations. First, the full symbolic model is split into background and perturbation systems. Cubic Hermite splines are then constructed for all unknown variables (like $a(\tau)$) to pass their solution from one stage to the next. This type of splines is perfect for interpolating ODEs, as cubic Hermite splines increase accuracy by taking both $\boldsymbol{u}(\tau)$ and $\boldsymbol{u}^{\prime}(\tau)$ into account, and $\boldsymbol{u}^{\prime}(\tau)$ is known analytically by definition (3). In contrast, observed variables (like $\Psi(\tau,k)$) are computed from the (splined) unknowns because their derivatives are not known analytically. Splining them directly would be less accurate. Modelers can simply write down a new background variable and access it in the perturbations for free. They can access any variable anywhere as if solving the entire set of equations at once, while still benefiting from splining under the hood.
@@
The separation into background and perturbation stages just reflects the perturbative structure of the linearized Einstein-Boltzmann equations, where each order depends only on those before it, and is always guaranteed. However, they can often be broken further down: most thermodynamics (recombination) models can be separated from the background, and some variables have integral solutions (e.g. the optical depth $\kappa(\tau)=\int_{\tau_{0}}^{\tau}\kappa^{\prime}(\bar{\tau})\mathrm{d}\bar{\tau}$ or line-of-sight integration) that can be computed in isolation after solving the differential equations. Over time, SymBoltz aims to extend the automatic background-perturbation separation by automatically incorporating more or all stages of the equations at hand. [^9] All stages can in principle be identified from a directed dependency graph between variables (e.g. from the Jacobian matrix). Cycles in the graph correspond to interdependent equations that must be integrated together in one ODE system (e.g. background and perturbations). Leaves with differentiated variables correspond to integral solutions. This feature is a major advantage with access to the symbolic equations.
@@ 2.1.4 Automatic solution interpolation
SymBoltz integrates equations in conformal time $\tau$ for several given wavenumbers $k$, and stores the result in an object that wraps independent ODE solutions for the background and perturbations for every $k$. Conveniently, this solution object can be queried for any variable or symbolic expression
$x(\tau,k)=x(\boldsymbol{u}(\tau,k_{i})).$ (4)
at any $\tau$ and $k$. The solution object automatically expands $x$ in terms of unknowns $\boldsymbol{u}$ (if $x$ is observed), and interpolates between solved wavenumbers $k_{i}$ (if $x$ is perturbative) and in $\tau$ using the ODE solver’s built-in interpolation method (or as in section 2.1.3 if $x$ is splined). This provides the modeler easy access to any variable and expression defined in the model. The solution interpolation is also incorporated into convenient plot recipes for trivial visualization of any expressions for any $\tau$ and $k$.
@@ 2.1.5 Automatic Jacobian generation and sparsity detection
Just like SymBoltz generates code for $\boldsymbol{f}$ in the ODE (3), it can use the same symbolic equations to generate its Jacobian
@@
Jacobians are crucial for solving stiff ODEs accurately and efficiently with implicit solvers (see section 2.2)! Manually coding them is an extremely tiresome and error-prone undertaking that must be repeated for model modifications. Evaluating them numerically with finite differences is approximate and expensive. Automation of this procedure is a powerful feature that again enables the modeler to focus solely on the main equations.
Another benefit is that analytical knowledge of the Jacobian lets one find its exact sparsity pattern. Numerical approaches to this are less robust and can mistake local zeros for global zeros. This permits constant factorization of sparse Jacobians, which improves performance of implicit solvers (see section 2.2). Full support for sparse Jacobians is still ongoing work, however.
SymBoltz can fall back from symbolic to automatic differentiation for the Jacobian. Finite differences can also be used.
2.1.6 Automatic change of variable (future work)
SymBoltz consistently formulates all differential equations with respect to conformal time $\tau$ as the independent variable. This is perhaps the most common parametrization in the literature. It is very natural because the equations are autonomous with respect to $\tau$ (i.e. $f(\tau,\boldsymbol{u})\rightarrow f(\boldsymbol{u})$, except some multipole truncation schemes that use $1/(k\tau)$, but these are unphysical).
Some other codes inconsistently use different independent variables in different stages. This requires manual translation into cosmic time $t$, redshift $z$, scale factor $a$ or its logarithm $b=\log a$, for example, involving differential transformations like $\mathrm{d}t=a(\tau)\,\mathrm{d}\tau$, $\mathrm{d}z=-(a^{\prime}(\tau)/a(\tau)^{2})\,\mathrm{d}\tau$, $\mathrm{d}a=a^{\prime}(\tau)\,\mathrm{d}\tau$ and $\mathrm{d}b=(a^{\prime}(\tau)/a(\tau))\,\mathrm{d}\tau$. This is another mechanical and error-prone process.
In the future, SymBoltz could automate change to another independent variable (which should be one-to-one with $\tau$) by transforming the symbolic expressions. This can be particularly useful to compute observables as a direct function of redshift, although one can invert $z(\tau)$. This is common in fits to observations.
2.1.7 Automatic unit handling (future work)
Most variables in SymBoltz are defined in internal dimensionless units (see appendix A). In the future, SymBoltz could use its symbolic pre-processing for more powerful unit features. For example, equations could be checked for dimensional validity to catch modeling mistakes. Input quantities could be automatically transformed from the user’s preferred (dimensionful) units (e.g. Mpc) to internal (dimensionless) units, and then back to the user’s units for output.
2.1.8 Automatic gauge transformation (future work)
SymBoltz is currently formulated only in the conformal Newtonian gauge. In the future, SymBoltz could conveniently transform the symbolic equations to other gauges automatically, such as the synchronous gauge used in CAMB and CLASS.
2.1.9 Automatic initial conditions (future work)
SymBoltz currently uses adiabatic initial conditions for the perturbations. In the future, SymBoltz could generate initial conditions for arbitrary models from the equations. This typically involves mechanical procedures like perturbation theory or power series expansion that could be automated.
2.1.10 Automatic approximation schemes (future work)
SymBoltz is initially designed to be free of approximation schemes (see section 2.2). In the future, SymBoltz could provide convenience utilities to help derive approximations to the equations. For example, it could automatically expand equations in power series of some smallness parameter. This could accelerate derivation of approximations for modified models to improve their performance with less effort.

### v2

@@ 2 Code architecture and main features
SymBoltz is designed around three main features. In short, it has a symbolic-numeric abstraction interface where users enter high-level symbolic equations, and automatically compiles them to fast low-level numerical functions. It cures stiffness with implicit ODE solvers, and uses no approximation schemes to keep models simple, elegant and extensible. It is differentiable, so one can get accurate derivatives of any output with respect to any input. This provides rapid model prototyping, helpful abstractions and automates tasks one must do manually when modifying other codes. SymBoltz encourages interactive use and pursues a modular design that lets users integrate what they need from the package into their own applications. The end goal is to prioritize the modeler, who should be able to just write down their equations while SymBoltz handles modeling chores.
The code is written in the Julia programming language (Bezanson et al., 2017), with a rich ecosystem of scientific packages and an aim to resolve the two-language problem. The symbolic-numeric interface is built on ModelingToolkit.jl (Ma et al., 2022) and Symbolics.jl (Gowda et al., 2022). The compiled functions are integrated by implicit ODE solvers in OrdinaryDiffEq.jl (Rackauckas and Nie, 2017). Automatic differentiation works through ForwardDiff.jl (Revels et al., 2016).
The next subsections describe SymBoltz’ three main features in more detail. Other implementation details are given in appendix A. This paper is as of SymBoltz version 0.14.1. Please see the package documentation for definitive up-to-date information.
@@ 2.1 Symbolic-numeric interface
The core of SymBoltz is built around a symbolic-numeric interface illustrated in fig. 1. Variables and equations are specified in a high-level user-friendly symbolic modeling language, then compiled down to low-level numerical code that is integrated by ODE solvers. With knowledge of the symbolic equations, SymBoltz can analyze their structure programmatically to ease model specification, automate related boilerplate tasks and generate fast and stable code that avoids approximations.
We think three properties motivate some kind of symbolic abstraction layer around the Einstein-Boltzmann equations.
First, the equations are most easily written as one large ODE, but optimally solved as several ODEs in stages like the background, perturbations and line-of-sight integration. Mathematically, the Jacobian matrix of this single ODE encodes which variables are independent and belong to which stages. It is easiest for modelers to specify equations in one common system, where the perturbations can refer to variables in the background, for example. The separation into computational stages can be automated.
Second, solving each stage is hard due to stiffness, performance requirements and its dependence on previous stages. To overcome this, it is helpful to automatically generate ODE code that is optimal in speed, accuracy and stability, generate the ODE Jacobian needed by implicit solvers and interpolate variables from previous stages. Modelers should not worry about this.
Third, the equations are often modified because the true cosmology is unknown. This can be made easier by organizing equations in a way that reflects the physical structure of the model, so it is easy to replace a related subset of equations with another without duplicating the entire model. This could be slow to do at the numerical level, but can be done with no runtime penalty in an intermediate symbolic representation of the equations.
@@
Inspired by this, SymBoltz inverts the traditional layout of Boltzmann solvers. While most codes are built as pipelines that follow the background, perturbations and other stages of the equations, SymBoltz is primarily structured around the physical components that they consists of. Related variables and equations are grouped in distinct components for the metric, gravitational theory, photons, baryons, dark matter, dark energy, neutrinos and other species (see fig. 1). This isolates everything related to one submodel to one place. Any set of such submodels are joined into a full cosmological model, like $\mathrm{\Lambda CDM}$. Interactions (e.g. Compton scattering or sourcing of gravity) are equations between components. Adding new submodels is easy, and SymBoltz is largely devoted to simply building a well-organized library of them.
SymBoltz then takes a full cosmological model, separates it into stages and generates numerical code to solve each stage. This preprocessing does not slow the code at runtime; to the contrary the symbolic equations give extra information like the Jacobian that enhances speed and stability. Modifications automatically enjoy these benefits without extra user input.
The modular component-based structure scales well in model space. For example, modified gravity or dark energy models can be written down as self-contained submodels and combined with other components to create many different extended cosmological models. The generic symbolic framework can also be used to build reduced models with non-interacting radiation, matter and dark energy species. When exploring modified gravity theories, it can be helpful to test such toy models to understand how gravity responds to pure fluids without coupling to baryons and photons.
In contrast, the monolithic layout of traditional codes scatter changes related to one submodel across different modules for input/output, the background, perturbations and so on. As more and more submodels are added, the code is increasingly intertwined and fragmented. They are also not flexible enough to test simpler toy models, as baryons and photons are hardcoded as fundamental to the structure of the code. This design fits the computational structure better than the physical one. It can result in an overwhelming code that becomes hard to read and modify, and whose complexity scales poorly in model space.
@@
As an alternative to component-based modeling, SymBoltz also includes a version of the default $\mathrm{\Lambda CDM}$ model with all equations packed into one big ?unstructured? system. This makes it trivial to change anything in the equations, but sacrifices modularity. The symbolic interface makes it very compact: SymBoltz defines a full $\mathrm{\Lambda CDM}$ model in just 277 straightforward lines of code [^8] https://hersle.github.io/SymBoltz.jl/stable/unstructured/, while the equivalent code to browse in CLASS is spread over 10 files with 27721 lines [^9] Counting input.{h,c}, background.{h,c}, thermodynamics.{h,c}, perturbations.{h,c} and wrap_recfast.{h,c} with wc -l..
PyCosmo’s symbolic interface also generates code for ODEs and sparse Jacobians, but does not focus on component-based modeling or automating tasks like stage separation, for example.
SymBoltz’ symbolic-numeric interface simplifies solving and modifying the Einstein-Boltzmann equations. It aims to maximize speed, stability and convenience with minimal user input. The next sections go into detail on how this is done. Section 3.2 shows a concrete example and some of its features in action.
@@ 2.1.1 Automatic numerical code generation
$\frac{\mathrm{d}\boldsymbol{u}}{\mathrm{d}\tau}=\boldsymbol{f}(\tau,\boldsymbol{u},\boldsymbol{p}).$ (3)
The generated code is fast and prevents users unfamiliar with Julia or SymBoltz from writing slow code. If necessary, one can escape the standard code generation and call arbitrary numerical functions, for example to solve a nonlinear equation for the minimum of a potential or interpolate tabulated data. The code generation handles tasks like allocating indices for each ODE state $u_{i}$, and performs optimizations such as common subexpression elimination (see section 3.2). This is helpful as Boltzmann solvers tend to have big $\boldsymbol{f}$ and ODE solvers call $\boldsymbol{f}$ many times.
@@ 2.1.2 Automatic handling of observed variables
A general ODE (3) has two types of variables: ?unknowns? $\boldsymbol{u}(\tau)$ are integrated with respect to time, while ?observed? variables are any functions of the unknowns. The Einstein-Boltzmann equations are usually written with many observed variables.
For example, consider the metric and gravity equations in sections A.1 and A.2 sourced by some arbitrary $\rho(\tau)$, $\delta\rho(\tau,k)$ and $\Pi(\tau,k)$. Here $a(\tau)$ and $\Phi(\tau,k)$ are the only unknown (differential) variables that the ODE is solved for, while one can ?observe? $z(\tau)$, $\mathscr{H}(\tau)$, $H(\tau)$ and $\Psi(\tau,k)$ from the unknowns. In addition, observed derivatives like $\mathscr{H}^{\prime}$ or $\Psi^{\prime}$ are not always trivial to express, but SymBoltz automatically expands them using the definitions of $\mathscr{H}$ or $\Psi$. Of course, one can eliminate all observeds by explicitly inserting them into the equations for the unknowns. But this reduces readability, as observed variables are helpful intermediate definitions that break up the equations, and one may want to extract them from the solution as well. Furthermore, modified models can change the sets of unknown and observed variables (e.g. modified gravity can change the constraint equation for $\Psi$ into a differential equation). It is easier to compose models when variables are not hardcoded as either unknown or observed.
SymBoltz reads in a full system of equations like that defined by appendix A and separates variables into unknowns and observeds. It then generates numerical code for solving the ODE for the unknowns (see section 2.1.1), but can automatically recompute any observed variables from the solution of the unknowns.
The bottom line is that users can easily use any variable anywhere by just referring to it, whether it is unknown or observed. In other solvers one may have to look up the expression for observed variables and recompute them manually. This is tedious, error-prone and can tempt users to make simplifying assumptions (e.g. approximating $\Psi\approx\Phi$ without anisotropic stress).
@@ 2.1.3 Automatic stage separation and splining of unknowns
In principle, one can solve the Einstein-Boltzmann equations by integrating the entire system (e.g. background and perturbations) at once. However, the system can be broken down into sequential computational stages that each depend only on those before it. To alleviate stiffness in each stage, separate concerns by solving every equation only once (e.g. avoid recomputing the background for every perturbation mode) and to improve performance, all Boltzmann codes solve the system stage-by-stage and spline variables from one stage as input to the next.
@@
To illustrate, again consider the general relativistic equations in section A.2 sourced by some known $\rho(\tau)$, $\delta\rho(\tau,k)$ and $\Pi(\tau,k)$. Clearly $\Phi(\tau,k)$ and $\Psi(\tau,k)$ depend on $a(\tau)$, but $a(\tau)$ does not depend on $\Phi(\tau,k)$ or $\Psi(\tau,k)$, reflecting the perturbative nature of the problem. One can first solve for only $a(\tau)$ in the ?background?, then spline and look up $a(\tau)$ to solve for $\Phi(\tau,k)$ and $\Psi(\tau,k)$ in the ?perturbations?, instead of solving for all three together and repeatedly solve for $a(\tau)$ for every $k$.
@@
SymBoltz uses the same stage separation strategy as any other Boltzmann code, but automates it with its knowledge of the symbolic equations. First, all equations are split into background and perturbation stages. Cubic Hermite splines are then constructed for all background unknowns (like $a(\tau)$) and passed on to the perturbations. This spline type is optimal for interpolating ODEs, as Hermite splines take both $\boldsymbol{u}(\tau)$ and $\boldsymbol{u}^{\prime}(\tau)$ into account for better accuracy, and $\boldsymbol{u}^{\prime}(\tau)$ is known analytically from $\boldsymbol{u}(\tau)$ and definition (3). In contrast, observed variables (like $\mathscr{H}(\tau)$) are computed from the (splined) unknowns because their derivatives are not known directly, and splining them is less accurate. This makes any background variable available in the perturbations ?for free?. All stages of the Einstein-Boltzmann equations are solved as if it is written as one combined system of equations, while splining and stage separation are done automatically under the hood.
@@
The separation into background and perturbation stages is guaranteed and just reflects the perturbative structure of the linearized Einstein-Boltzmann equations, where each order depends only on those before it. However, they can often be broken further down: most thermodynamics (recombination) models can be separated from the background, and some variables have integral solutions (e.g. the optical depth $\kappa(\tau)=\int_{\tau_{0}}^{\tau}\kappa^{\prime}(\bar{\tau})\mathrm{d}\bar{\tau}$ or line-of-sight integration) that can be computed in isolation after solving the differential equations. In the future, SymBoltz’ background-perturbations separation could be extended to automatically split all smaller stages by symbolically inspecting the dependencies between variables in the equations.
@@ 2.1.4 Automatic solution interpolation
SymBoltz integrates the background and several perturbation ODEs in conformal time $\tau$ for several wavenumbers $k$. All the results are stored in a special solution object. This solution object can be queried for any variable or symbolic expression
$S(\tau)\quad\text{or}\quad S(\tau,k).$ (4)
This looks up the background solution if $S$ is a background variable, or the perturbation solutions if it is perturbative. It interpolates from the times stored by the ODE solver to the requested $\tau$ using the solver’s dense interpolation. If $S$ is perturbative, it also interpolates from the solved perturbation $k$-modes to the requested $k$. If $S$ is an unknown, it is returned directly from the ODE solution. If $S$ is observed, it is instead recomputed automatically from the equation that defines it as a function of the unknowns.
The result is that the user can access any variable in the model without having to do this manually. The solution interpolation is also incorporated into plot recipes for visualization of any variable as a function of $\tau$ and/or $k$.
@@ 2.1.5 Automatic Jacobian generation and sparsity detection
Just like SymBoltz generates code for $\boldsymbol{f}$ in the ODE (3), it uses the same symbolic equations to generate its Jacobian $\boldsymbol{J}$ with entries
@@
Jacobians are crucial for solving stiff ODEs with implicit solvers. Manually coding them is tiresome, error-prone, and must be repeated for new models. Numerical evaluation with finite differences is both approximate and slow. It can also be computed with automatic differentiation, but we find this slower and overkill compared to the analytical $\boldsymbol{J}$. Analytical Jacobians are as fast and stable as possible, and automatic generation lets the solver benefit from it while the modeler focuses on the high-level equations.
SymBoltz stores the Jacobian in sparse form to boost performance when it has many zeros. From the analytical $\boldsymbol{J}$, SymBoltz precomputes its exact sparsity pattern (where $J_{ij}=0$). This is hard to specify manually, and numerical solvers cannot distinguish false local zeros (for some input) from true global zeros (for all inputs). Without approximations, SymBoltz then reuses the same sparsity structure throughout the integration, computes the nonzero $J_{ij}$ analytically and stores them directly in the sparse $\boldsymbol{J}$ with no extra cost from looking up sparse matrix indices.
Easy, accurate, fast and sparse Jacobians are crucial for SymBoltz to remain performant without approximations! It is a major advantage with the symbolic approach, as in PyCosmo.